# Predicting Change, Not States: An Alternate Framework for Neural PDE Surrogates

### v1

@@ 3.3 Equations Considered
where we uniformly sample $A_{j}\in[-0.5,0.5],\>\omega_{j}\in[-0.4,0.4],\>l_{j}\in\{1,2,3\},\>\phi_{j}\in[0,2\pi)$ while fixing $J=5,L=16$. Samples are generated at a spatial resolution of $n_{x}=100$ on a domain $x\in[0,16]$. The Advection and Heat equations are solved from $t=0$ to $t=2$, while the KS equations are solved from $t=0$ to $t=100$ due to the different dynamics. Additionally equations coefficients are uniformly sampled for the Advection and Heat equations to introduce additional difficulty: $c\in[0.1,2.5],\>\nu\in[0.1,0.8]$. For the KS equation, coefficients are kept fixed to the more challenging dynamics. Periodic boundary conditions are used for all equations. Lastly, 4096 training samples and 256 validation samples are used for all PDEs.
While conventionally a minor detail, we find that care must be taken when choosing the temporal discretization of the dataset. In state prediction, the temporal resolution must be fine enough to resolve important dynamics, yet not too fine to introduce excessive error propagation during inference. For many PDEs, such as turbulence, this can be a compromise where one objective cannot be improved without sacrificing the other [48]. In the context of derivative prediction, the temporal discretization is less restrictive since the model is not constrained to a resolution defined by the dataset. However, during inference, the timestep cannot be excessively large as to introduce error from the ODE integrator and shift the input distribution to the model. For the main experiments, the discretization is chosen to be 125 timesteps ($\Delta t=0.016s$) for the Advection and Heat equations, and 400 timesteps ($\Delta t=0.25$) for the KS equation, although these will be varied when investigating the effect of varying timescales on neural surrogate performance.
@@ 4.1 Prediction Accuracy
In all experiments, we consider evaluating the method using a Fourier Neural Operator (FNO) [3] and Unet [10], which are two popular model choices for benchmarking uniform-grid PDE prediction problems [49, 50]. Models are either trained with the state prediction objective (State Pred.), or the derivative prediction objective with a 4-th order Richardson’s extrapolation to ensure the accuracy of the training labels. Furthermore, different integration schemes are evaluated during inference. Models are evaluated on a validation set by computing either the rollout error or correlation time. Rollout error is defined to be the relative L2 error of the predicted, auto-regressive trajectory with respect to the true trajectory, averaged over all timesteps. In chaotic systems such as the KS equation, rollout error can be heavily skewed once model predictions diverge, therefore correlation time is used to judge model performance. Correlation time is defined to be the timestep after which model predictions have a Pearson correlation of less than 0.8 with respect to the ground truth.
PDE: Adv Heat KS Burgers NS
Metric: Roll. Err. $\downarrow$ Roll. Err. $\downarrow$ Corr. Time $\uparrow$ Roll. Err. $\downarrow$ Roll. Err. $\downarrow$
FNO (State Pred.) 0.544 0.589 140.75 0.437 0.715
FNO (FwdEuler) 0.048 0.141 85.75 0.196 0.159
FNO (Adams) 0.032 0.141 177 0.174 0.100
FNO (Heun) 0.032 0.141 183 0.175 0.100
FNO (RK4) 0.033 0.141 184 0.175 0.100
Unet (State Pred.) 0.386 0.199 70 Unstable Unstable
Unet (FwdEuler) 0.125 0.141 65.75 0.444 0.384
Unet (Adams) 0.117 0.141 79.5 0.397 0.334
Unet (Heun) 0.117 0.141 79.75 0.400 0.333
Unet (RK4) 0.116 0.141 80.25 0.402 0.333
Table 1: Results on prediction accuracy across different PDEs, models, and training/inference frameworks.
@@
Validation metrics are reported in Table 1. We find that derivative prediction can achieve much lower errors than state prediction across equations and model choices. One hypothesized reason for this is because derivative prediction can be more stable than state prediction; both models are trained such that the one-step training error is very low, however, during inference, using state prediction leads to error accumulation and even unstable rollouts. From this point of view, incorporating an ODE integrator during inference can help stabilize neural surrogate predictions by using a numerical, physics-based prior to update the solution.
@@
For equations with simple dynamics, such as the Advection, Heat, and 2D Burgers equations, we observe that using higher-order ODE integrators does not significantly affect accuracies, although the Forward Euler method can perform worse. This doesn’t mean higher-order integrators aren’t useful; they can still accelerate inference times by taking larger timesteps during inference due to being more robust. However, within this experimental setup, the Adams-Bashforth method can be a good choice during inference due to its higher accuracy and identical inference cost to the Forward Euler method or state prediction. For more challenging systems, such as the KS equation, higher-order integrators can be more accurate, since coarser integrators can accumulate larger errors and shift the input distribution during inference. Lastly, we note that the ODE integrator can be changed during inference without re-training the model, allowing the speed or accuracy of models trained with derivative prediction to be easily changed.
Additionally, we plot sample rollouts of FNO and Unet models trained using state prediction or derivative prediction with an RK4 integrator. Results in 1D and 2D are plotted in Figure 2 and Figure 3. Interestingly, FNO models struggle with simple 1D dynamics, which is corroborated by the validation metrics. We believe that this is due to overfitting, where FNO models largely optimize for one-step accuracy without long-term stability. Indeed, when applying pushforward or refiner methods in the following section, the error of FNO methods on simple 1D problems improves drastically. Unet models perform better on the Advection and Heat equations; however, stability is still a challenge, with spurious predictions being made at later timesteps. Applying derivative prediction can fix these issues and lead to stable, accurate predictions for simple 1D PDEs. For more chaotic 1D systems, such as the KS equation, derivative prediction is able to delay error accumulation and extend the correlation time of predictions. Furthermore, derivative prediction can stabilize Unet predictions, despite being inaccurate in the long term.
For 2D systems, derivative prediction can similarly stabilize FNO and Unet predictions. Unet models are especially prone to error propagation, since FNO models have helpful inductive biases that create sinusoidal outputs. Applying derivative prediction can result in accurate solutions; however, for more complex systems (NS), solutions eventually lose accuracy.
@@
Figure 2: Comparison of state prediction and derivative prediction using an RK4 integrator on 1D PDEs. Time is plotted on the x-axis and nodal values on the y-axis. Correlation time is denoted for the KS equation.
Figure 3: Comparison of state prediction and derivative prediction using an RK4 integrator on 2D PDEs. The spatial dimensions are plotted for each frame, at multiple snapshots in time from top to bottom.
@@ 4.2 Comparison to Training Modifiers
We train models using various modifications with the state prediction framework and additionally report baseline models using only state prediction and derivative prediction with an RK4 integrator. Validation metrics are reported in Table 2. We reproduce previous results showing the effectiveness of scaling model size, unrolled training, and refiner methods, which tend to improve the performance of neural surrogates across many PDEs and models. However, we observe the improvement of these modifications can be more nuanced. For example, scaling the model size can depend on the architecture’s scalability and capacity, as well as data constraints to prevent overfitting. Indeed, we observe that FNO models tend to overfit in simple test cases (Advection, Heat), while Unet models do not, underlining previous results on the scalability of Unet vs. FNO models in PDE contexts [16]. Furthermore, while refiner methods are very powerful, in some cases achieving the best performance, introducing a denoising objective creates additional complexity during training, and in rare cases performs worse than a baseline. Refiner methods can also be dependent on model architecture, with some models performing better than others on denoising; in particular, FNO models can struggle with spurious high frequencies introduced by Gaussian noise that are exacerbated in 2D PDEs.
Lastly, using the pushforward trick tends to consistently improve model performance, likely due to the simplicity of the method in improving robustness to model noise. Likewise, the simplicity of using a different training objective in derivative prediction also leads to consistent improvements in model performance, outperforming unrolled training and suggesting that derivative prediction can be generally applicable across different architectures and problems. In addition, derivative prediction does not add additional training cost, while other training modifications generally require more compute and memory.
@@
It is also interesting to note that these modifications are not exclusive of each other; in fact, it is straightforward to implement derivative prediction in conjunction with scaling the model size, using unrolled training, or applying refiner methods to further improve performance. As a final remark, while overall performance gains of these methods support prior work and observations, applying these training modifications successfully still requires an understanding of the problem setup, model capacity, dataset size, or relevant theory, as the results show that training intricacies and exceptions to overall trends still exist. Much like any numerical or machine-learning method, using derivative prediction similarly has these caveats, however, we leave this discussion to later sections.
PDE: Adv Heat KS Burgers NS
Metric: Roll. Err. $\downarrow$ Roll. Err. $\downarrow$ Corr. Time $\uparrow$ Roll. Err. $\downarrow$ Roll. Err. $\downarrow$
FNO (State Pred.) 0.544 0.589 140.75 0.437 0.715
FNO (4x params) 0.826 0.991 156.25 0.276 0.599
FNO (Pushforward) 0.357 0.486 146 0.397 0.090
FNO (Refiner) 0.036 0.167 148.75 1.141 0.596
FNO (RK4) 0.033 0.141 184 0.175 0.100
Unet (State Pred.) 0.386 0.199 70 Unstable Unstable
Unet (4x params) 0.370 0.147 67.25 Unstable 0.457
Unet (Pushforward) 0.276 0.169 76.75 Unstable Unstable
Unet (Refiner) 0.275 0.175 63.5 0.313 0.235
Unet (RK4) 0.116 0.141 80.25 0.402 0.333
Table 2: Comparing different training modifications on various PDEs and models.
4.3 Inference Modifications
@@
FNO (FwdEuler) 0.048 0.159 Unet (FwdEuler) 0.125 0.384
FNO (FwdEuler + 2x data) 0.045 0.139 Unet (FwdEuler + 2x data) 0.119 0.345
FNO (FwdEuler + 2x steps) 0.037 0.120 Unet (FwdEuler + 2x steps) 0.120 0.346
FNO (Heun) 0.033 0.100 Unet (Heun) 0.117 0.333
Table 3: Comparing different inference modifications across various models and PDEs.
@@
We report the results of training on more finely discretized data (2x data) and taking smaller timesteps during inference (2x steps) in Table 3. Results are evaluated for the Advection and Navier-Stokes equations to validate these ideas on a simple 1D case and a more complex 2D case. In addition, we also report baselines using the Forward Euler or Heun’s method without modifications. Interestingly, training on more finely discretized data can result in lower rollout errors without additional inference cost, likely from learning more accurate derivative estimates. This effect is also more apparent in more complex systems (NS), where more data and better derivative estimates become more important. This opens opportunities for models trained using derivative prediction to discard less data than conventional state prediction frameworks and achieve better performance. However, we hypothesize this can have diminishing returns, as further increasing the resolution of the dataset results in adding data samples that are already very similar to the existing dataset.
@@ 4.4 Prediction Timescales
(a) 1D Advection equation.
(b) 2D Navier-Stokes equation.
Figure 4: Rollout error of FNO or Unet models trained with state prediction or derivative prediction with a Forward Euler or RK4 integrator; errors are plotted after predicting solutions at different timescales $\Delta t$. The resolution where $CFL=1$ is denoted for 1D Advection. Empty circles denote unstable rollouts.
@@ 5 Discussion
Figure 5: Visualization of the loss landscape of state prediction and derivative prediction after training a Unet on 1D Advection. We find that the derivative prediction objective can be empirically easier to learn.

### v2

@@ 3.2 Inference
3.3 Models Considered
We consider evaluating the training and inference frameworks using the Fourier Neural Operator (FNO) [3] and Unet [13], which are two popular model choices for benchmarking uniform-grid PDE prediction problems [63, 64]. FNO models are a widely used neural operator architecture that approximates an infinite-dimensional mapping between input and output Banach spaces. This is accomplished by using the integral kernel operator $u(x)=\int_{D}\kappa(x,y)v(y)\text{d}y$ to map between from an input function $v(y)$ to an ouput function $u(x)$. Learning the kernel $\kappa_{\theta}(x,y)$ in this operator is a convenient way to parameterize an infinite-dimensional mapping; however in practice, the full integral kernel operator is expensive to both train and evaluate. While several works have seeked to mitigate this computational cost [65, 66, 67], the insight behind FNO models is that the integral kernel operator can be expressed as the product of the Fourier transforms of the kernel and the input function by the convolution theorem. Truncating the kernel and input Fourier series and using a Fast Fourier Transform allows the operator to be learned and evaluated quickly. In 1D, we use FNO models with $\sim 1M$ parameters, and in 2D, we use FNO models with $\sim 7M$ parameters. We refer readers to the provided implementation for additional details on network width, depth, and number of Fourier modes.
The theory behind FNO and neural operators as a whole guarantees an infinite-dimensional mapping [5]; however, in practice, models relying on convolutions, such as the Unet, have shown good performance despite not satisfying discretization invariance. As such, in addition to FNO models, Unet models are a common architecture choice due to being well-studied and optimized for uniform-grid prediction tasks [13]. Unets are widely used in the broader literature on computer vision and generative modeling [68, 69], allowing many of the advances in CNN architecture design to inform PDE prediction tasks. In particular, attention blocks are placed after convolutional blocks and Wide ResNet layers [70] are used. In 1D, we use Unet models with $\sim 3M$ parameters and in 2D, we use Unet models with $\sim 9M$ parameters. Additional details are given in the implementation for the network width, depth, and up/downsampling layers.
@@ 3.4 Equations Considered
where we uniformly sample $A_{j}\in[-0.5,0.5],\>\omega_{j}\in[-0.4,0.4],\>l_{j}\in\{1,2,3\},\>\phi_{j}\in[0,2\pi)$ while fixing $J=5,L=16$. Samples are generated at a spatial resolution of $n_{x}=100$ on a domain $x\in[0,16]$. The Advection and Heat equations are solved from $t=0$ to $t=2$, while the KS equations are solved from $t=0$ to $t=100$ due to the different dynamics. The first 50 seconds of the KS equation are discarded to allow the system burn-in and to converge to the chaotic state. Additionally equations coefficients are uniformly sampled for the Advection and Heat equations to introduce additional difficulty: $c\in[0.1,2.5],\>\nu\in[0.1,0.8]$. For the KS equation, coefficients are kept fixed to the more challenging dynamics. Periodic boundary conditions are used for all equations. Lastly, 4096 training samples and 256 validation samples are used for all PDEs.
While conventionally a minor detail, we find that care must be taken when choosing the temporal discretization of the dataset. In state prediction, the temporal resolution must be fine enough to resolve important dynamics, yet not too fine to introduce excessive error propagation during inference. For many PDEs, such as turbulence, this can be a compromise where one objective cannot be improved without sacrificing the other [71]; small timescales are need to resolve the energy cascade yet this results in high autoregressive error propagation. In the context of derivative prediction, the temporal discretization is less restrictive since the model is not constrained to a resolution defined by the dataset. However, during inference, the timestep cannot be excessively large as to introduce error from the ODE integrator and shift the input distribution to the model. For the main experiments, the discretization is chosen to be 125 timesteps ($\Delta t=0.016s$) for the Advection and Heat equations, and 200 timesteps ($\Delta t=0.25$) for the KS equation, although these will be varied when investigating the effect of different timescales on neural surrogate performance.
@@ 4 Results
PDE: Adv Heat KS Burgers NS Kolm. Flow
Metric: Roll. Err. $\downarrow$ Roll. Err. $\downarrow$ Corr. Time $\uparrow$ Roll. Err. $\downarrow$ Roll. Err. $\downarrow$ Corr. Time $\uparrow$
FNO (State Pred.) 0.498 0.589 139.75 0.437 0.715 51.4
FNO (FwdEuler) 0.048 0.141 79.25 0.196 0.159 29.0
FNO (Adams) 0.032 0.141 183 0.174 0.100 45.5
FNO (Heun) 0.032 0.141 197.75 0.175 0.100 81.6
FNO (RK4) 0.033 0.141 198 0.175 0.100 82.3
Unet (State Pred.) 0.044 0.149 153.5 0.666 0.099 35.1
Unet (FwdEuler) 0.032 0.139 76.5 0.280 0.093 23.3
Unet (Adams) 0.011 0.139 167.75 0.264 0.048 27.9
Unet (Heun) 0.010 0.139 173.5 0.263 0.049 53.7
Unet (RK4) 0.010 0.139 174.5 0.264 0.049 88.9
Table 1: Prediction Accuracy. Results on prediction accuracy across different PDEs, models, and training/inference frameworks.
@@ 4.1 Prediction Accuracy
To compare the prediction accuracy of different training frameworks, FNO or Unet models are either trained with the state prediction objective (State Pred.), or the derivative prediction objective, where labels are calculated using a 4-th order Richardson’s extrapolation. Furthermore, different integration schemes are compared during inference. Models are evaluated on the validation set by computing either the rollout error or correlation time. Rollout error is defined to be the relative L2 error of the predicted, auto-regressive trajectory $\hat{\mathbf{u}}$ with respect to the true trajectory $\mathbf{u}$, averaged over all timesteps, which is given by: $\frac{1}{T}\sum_{i=1}^{T}\frac{||\hat{\mathbf{u}}(t_{i})-\mathbf{u}(t_{i})||^{2}}{||\mathbf{u}(t_{i})||^{2}}$. In chaotic systems such as the KS equation or Kolmogorov Flow, rollout error can be heavily skewed and go to infinity once model predictions diverge, therefore correlation time is used to judge model performance. Correlation time is defined to be the timestep after which model predictions have a Pearson correlation of less than 0.8 with respect to the ground truth.
@@
Validation metrics are reported in Table 1. We find that derivative prediction can achieve lower errors than state prediction across equations and model choices. One hypothesized reason for this is that derivative prediction can be more stable than state prediction; both models are trained such that the one-step training error is low, however, during inference, using state prediction leads to error accumulation and even unstable rollouts. Since derivative prediction simply adds a change to the previous state rather than constructing a new state, it may be able to mitigate error accumulation if the previous state is assumed to be accurate, or the timestep $\Delta t$ is small enough such that the state change is small.
Figure 2: 1D KS Equation. Comparison of state prediction and derivative prediction using an RK4 integrator on the 1D KS equation. Time is plotted on the x-axis and nodal values on the y-axis. Correlation time is denoted with a dashed white line.
Figure 3: 2D Kolmogorov Flow. Comparison of state prediction and derivative prediction using an RK4 integrator on 2D Kolmogorov Flow. The spatial dimensions are plotted for each frame, at multiple snapshots in time from top to bottom. The rollout is visualized at times $0,\frac{T}{3},\frac{2T}{3},T$, where T is the prediction horizon.
@@
For equations with simple dynamics, such as the Advection, Heat, and 2D Burgers equations, we observe that using higher-order ODE integrators does not significantly affect accuracies, although the Forward Euler method can perform worse. This doesn’t mean higher-order integrators aren’t useful; they can still accelerate inference times by taking larger timesteps during inference due to being more robust. However, at this resolution, the Adams-Bashforth method can be a good choice due to its higher accuracy and identical inference cost to the Forward Euler method and state prediction, due to relying on cached, past predictions. For more challenging systems, such as the KS equation or Kolmogorov flow, higher-order integrators can be more accurate, since coarser integrators can accumulate larger errors and shift the input distribution during inference. Lastly, we note that the ODE integrator can be changed during inference without re-training the model, allowing the speed or accuracy of models trained with derivative prediction to be easily changed.
Additionally, we plot sample rollouts of FNO and Unet models trained using state prediction or derivative prediction with an RK4 integrator. Results for the most challenging PDEs (KS, Kolm. Flow) are plotted in Figure 2 and Figure 3. Qualitatively, models using derivative prediction on chaotic and complex PDE systems tend to remain accurate for longer. This can be seen directly for the KS equation, and for the Kolmogorov flow experiment, the predicted states tend to diverge at later timesteps for derivative prediction models despite all models eventually diverging.
@@
PDE: Adv Heat KS Burgers NS Kolm. Flow
Metric: Roll. Err. $\downarrow$ Roll. Err. $\downarrow$ Corr. Time $\uparrow$ Roll. Err. $\downarrow$ Roll. Err. $\downarrow$ Corr. Time $\uparrow$
FNO (State Pred.) 0.498 0.589 139.75 0.437 0.715 51.4
FNO (4x params) 0.826 0.991 147.75 0.276 0.599 67.4
FNO (Pushforward) 0.357 0.486 141 0.397 0.090 53.3
FNO (Refiner) 0.036 0.167 159.75 1.141 0.596 37.5
FNO (RK4) 0.033 0.141 197.25 0.175 0.100 82.3
FNO (RK4+Pushforward) 0.023 0.140 194.25 0.322 0.058 83.8
Unet (State Pred.) 0.044 0.149 153.5 0.666 0.099 35.1
Unet (4x params) 0.036 0.144 145.25 0.222 0.053 52.1
Unet (Pushforward) 0.048 0.149 145 0.746 0.078 36.0
Unet (Refiner) 0.082 0.176 145.5 0.217 0.180 51.1
Unet (RK4) 0.010 0.139 174.5 0.264 0.049 88.9
Unet (RK4+Pushforward) 0.010 0.139 177 0.264 0.034 90.6
Table 2: Training Modifiers. Comparing different training modifications on various PDEs, models and frameworks.
@@ 4.2 Comparison to Training Modifiers
We train models using various modifications with the state prediction framework and additionally report baseline models using only state prediction and derivative prediction with an RK4 integrator. Validation metrics are reported in Table 2. We reproduce previous results showing the effectiveness of scaling model size, unrolled training, and refiner methods, which tend to improve the performance of neural surrogates across many PDEs and models. However, we observe the improvement of these modifications can be more nuanced. For example, scaling the model size can depend on the architecture’s scalability and capacity, as well as data constraints to prevent overfitting. Indeed, we observe that FNO models tend to overfit in simple test cases (Advection, Heat), while Unet models do not, underlining previous results on the scalability of Unet vs. FNO models in PDE contexts [24]. Furthermore, while refiner methods are very powerful, in some cases achieving the best performance, introducing a denoising objective creates additional complexity during training, and in certain cases performs worse than a baseline. Refiner methods can also be dependent on model architecture, with some models performing better than others on denoising; in particular, FNO models can struggle with spurious high frequencies introduced by Gaussian noise that are exacerbated in 2D PDEs. As a whole, derivative prediction can be a useful tool to improve model performance and can have improvements comparable or better than other training modifications. In addition, derivative prediction does not add additional training cost, while other training modifications generally require more compute and memory during training.
It is also interesting to note that these modifications are not exclusive of each other; in fact, it is straightforward to implement derivative prediction in conjunction with scaling the model size, using unrolled training, or applying refiner methods to further improve performance. We demonstrate this by training a model using derivative prediction (RK4) with the pushforward trick; the model makes a prediction of the current derivative $F_{\theta}(\mathbf{u}(t))=\frac{d\mathbf{u}}{dt}|_{t=t}$, it is integrated to obtain the next predicted solution $\hat{\mathbf{u}}(t+1)$, and this is used to predict the following derivative $F_{\theta}(\hat{\mathbf{u}}(t+1))=\frac{d\mathbf{u}}{dt}|_{t=t+1}$, which is finally used in the loss function. In simple cases, we find that the pushforward trick is unnecessary, however for more complex PDEs, combining this modification with derivative prediction can have modest improvements.
@@
As a final remark, while overall performance gains of these methods support prior work and observations, applying these training modifications successfully still requires an understanding of the problem setup, model capacity, dataset size, or relevant theory, as the results show that training intricacies and exceptions to overall trends still exist. Much like any numerical or machine-learning method, using derivative prediction similarly has these considerations, however, we leave this discussion to later sections.
@@
FNO (FwdEuler) 0.048 0.159 Unet (FwdEuler) 0.032 0.093
FNO (FwdEuler + 2x data) 0.045 0.139 Unet (FwdEuler + 2x data) 0.033 0.091
FNO (FwdEuler + 2x steps) 0.037 0.120 Unet (FwdEuler + 2x steps) 0.019 0.059
FNO (Heun) 0.033 0.100 Unet (Heun) 0.010 0.049
Table 3: Inference Modifiers. Comparing different inference modifications across various models and PDEs.
4.3 Inference Modifications
@@
We report the results of training on more finely discretized data (2x data) and taking smaller timesteps during inference (2x steps) in Table 3. Results are evaluated for the Advection and Navier-Stokes equations to validate these ideas on a simple 1D case and a more complex 2D case. In addition, we also report baselines using the Forward Euler or Heun’s method without modifications. Interestingly, training on more finely discretized data can result in lower rollout errors without additional inference cost, likely from learning more accurate derivative estimates. For simple systems (Adv), this additional data does not seem to be necessary, but for more complex systems (NS), more data and better derivative estimates become more beneficial. This opens opportunities for models trained using derivative prediction to discard less data than conventional state prediction frameworks and achieve better performance. However, we hypothesize this can have diminishing returns, as further increasing the resolution of the dataset results in adding data samples that are already very similar to the existing dataset.
@@
(a) 1D Advection equation.
(b) 2D Navier-Stokes equation.
Figure 4: Prediction Timescales. Rollout error of FNO or Unet models trained with state prediction or derivative prediction with a Forward Euler or RK4 integrator; errors are plotted after predicting solutions at different timescales $\Delta t$. The resolution at which $CFL=1$ is denoted for 1D Advection.
@@ 4.4 Prediction Timescales
Figure 5: Loss Landscapes. Visualization of the loss landscape of state prediction and derivative prediction after training a Unet on 1D Advection. Trained model parameters $\theta^{*}$ are varied by a linear combination of random direction vectors $\delta,\gamma$ at sampled $a,b$ values. Validation rollout loss for each perturbed set of weights $\theta^{*}+a\delta+b\gamma$ are calculated and averaged across the validation set, which is plotted as contours or surfaces. The scale and threshold of $L$ is kept consistent across contour and surface plots for a given model/framework.
5 Discussion